You manage a coffee shop. You begin the day with 25 pounds of coffee beans. You use 4 pounds of beans to brew coffee and sell 20 equal-weight bags of beans. At the end of the day you have 6 pounds of beans left. What is the weight of each bag of beans you sell?
Answer by edjones(8007)

25-4=21 lbs remain
21-6=15 lbs got sold (6 didn't)
15/20 = 3/4 lb weight of each bag sold\
.
Ed